### v1

@@ Abstract
Foundation model Kidney Pathology Instance Segmentation Nuclei Model Agreement Ensemble learning.
@@ 1 INTRODUCTION
To answer the preceding question, we evaluated the performance of current cell nuclei foundation models in kidney pathology. As shown in Fig. 1, we created a diverse evaluation dataset that includes kidney nuclei data from 2,542 kidney whole slide images (WSIs) sourced from humans and rodents across both public and in-house datasets. The dataset includes various tissue types and sizes, such as biopsy, autopsy, and nephrectomy samples, and features both formalin-fixed paraffin-embedded (FFPE) and frozen tissues. To our knowledge, the scale of this study’s kidney WSIs surpasses all publicly available labeled nuclei datasets that include the kidney, as illustrated in Fig. 1a. This work conducts a comparative analysis of three widely used state-of-the-art (SOTA) cell nuclei foundation models (FM) — Cellpose[14], StarDist [15], and CellViT [8] — on the task of nuclei instance segmentation. For validation, we include a manual rating-based system to curate foundation model predictions and identify failure samples across different models and modalities. To sum up, this study provides a comprehensive assessment of cell nuclei AI foundation models in the field of kidney pathology. The contribution of this paper is three-fold:
$\bullet$ Creation of a Diverse, Large-Scale Kidney Nuclei Dataset: The kidney dataset for evaluation consists of 2,542 kidney whole slide images (WSIs) collected from both human and rodent sources, encompassing various tissue types, sizes, and staining methods. To our knowledge, this is the largest-scale evaluation of its kind to date.
$\bullet$ Rating-based segmentation curation with Foundation Models: We utilized three state-of-the-art foundation models (Cellpose[14], StarDist [15], and CellViT [8]) in conjunction with a rating-based system to curate foundation model predictions. Then the evaluation can be shaped as analyzing distribution of the model predictions across the dataset.
$\bullet$ Agreement Analysis between Foundation Models: We performed agreement analysis to identify image patches where all models consistently rated them as ”bad.” It allowed us to summarize examples and patterns of these consensus ”failure” image patches, offering valuable insights for addressing nuclei segmentation challenges in kidney pathology.
Figure 1: Overall framework. (1) The creation of a diverse, large-scale kidney nuclei dataset.(2) Nuclei instance segmentation with Cellpose [14], StarDist [15], and CellViT [8]. (3) Rating on Segmentation Performance. Meaningful image patches were then identified through model agreement analysis.
2 Methods
Fig. 1 illustrates the overall framework of this study: (1) Large-scale diverse dataset creation, (2) Nuclei instance segmentation, (3) Rating on Segmentation Performance. Meaningful image patches were then identified through model agreement analysis.
@@ 2.2 Nuclei instance segmentation
Nuclei instance segmentation was performed on these kidney nuclei image patches using three foundation models: Cellpose [14], StarDist [15], and CellViT [8]. We summarized the three foundation models and provided GitHub links to our customized GPU-optimized inference pipelines in Table. 1:
@@
Year-Month Model Backbone Post-processing Customized Inference
2022 Mar StarDist[15] U-Net Star-convex Polygon GitHub Link
2022 Nov Cellpose[14] U-Net GradientFlow Tracking GitHub Link
2023 Oct CellViT[8] ViT HoVer-Net[6] GitHub Link
@@ 2.3 Rating on Segmentation Performance
Two pathologist-trained students scored the model predictions as “good,” “medium,” or “bad” based on a standard provided by a renal pathologist. “good” predictions correctly capture about 90% of nuclei in an image patch, “bad” predictions capture less than 50%, and the rest are categorized as “medium.” We apply this scoring system to quantitatively assess each model’s predictions in our dataset. Finally, we ensemble ratings from three foundation models to curate strongly agreed-upon failure samples. These image patches can be used for training or fine-tuning future foundation models in kidney pathology.
@@ 3 Experiments
3.1 Distribution of foundation model predictions
We analyzed the distribution of predictions for each foundation model across our evaluation dataset. This analysis provides insights into the performance and behavior of each model by examining the frequency of different prediction classes (”good”, ”medium”, ”bad”) within each foundation model. The prediction class assignments were determined through the manual rating-based curation.
@@ 3.2 Agreement analysis between foundation models
3.3 Consensus ”bad” image patches mining
@@ 4 Results
From Fig. 2, we can observe the distribution of predictions for each model. For all three foundation models, at least 40% of predictions are categorized as ”bad” or ”medium.” at least 40% of predictions are ”good”. This indicates that, despite not being specifically trained on kidney pathology nuclei data, the models retain some transferable knowledge applicable to our task. However, the 40% gap underscores the need for a specialized foundation model for kidney pathology. Cellpose and StarDist have similar amount of ”good” predictions, while CellViT shows a notably higher number of ”good” predictions of 5,609, which suggests its potentially superior performance and adaptability in kidney pathology.
@@
Fig. 3a illustrates the consistency between model predictions. As shown, no pair exhibits over 90% agreement or complete disagreement. The highest agreement is observed between Cellpose and StarDist (0.76), while the lowest is between Cellpose and CellViT (0.67). These non-extreme values, with a maximum of 0.76 and a minimum of 0.67, indicate that the current SOTA nuclei foundation models in digital pathology do not generalize exceptionally well to diverse large-scale kidney data. Each model has its own strengths, and this inconsistency inspires us to combine multiple SOTA foundation models to enhance performance in downstream kidney nuclei tasks.
Table 2: This table shows the percentages of model agreement in each prediction class
@@
Both Fig. 3b and Table. 2 show the number of image patches where all three models agree, two models agree, or no models agree for each prediction class (”good”, ”medium”, ”bad”). For better visualization, these numbers are normalized as percentages within each prediction class. It is evident that ”bad” cases exhibit the highest inter-model reliability, with over 70% agreement among all three models. Particularly, samples where all three foundation models agree in both the ”good” (Consensus ”good”) and ”bad” (Consensus ”failure”) classes are displayed, along with their respective failure categories. As shown, current nuclei foundation models perform well on image patches with darker nuclei and higher contrast between nuclei and non-nuclei regions. Conversely, Consensus ”failure” samples exhibit overall lower staining intensity, lighter nuclei, and reduced imaging contrast. From the failure categories in Fig. 3, it is evident that certain types of nuclei and tissue conditions diminish the segmentation performance of foundation models. Specifically, long and flat nuclei, nuclei with blurred boundaries, and densely distributed nuclei within glomeruli are particularly difficult to segment accurately. Additionally, slides that are too thin, contain fatty tissues, or have an excessive number of red blood cells can lead to a higher incidence of false positive predictions.
@@ 5 Conclusion
This study provides a comprehensive assessment of cell nuclei foundation models (Cellpose, StarDist, CellViT) in kidney pathology, focusing on nuclei instance segmentation. By evaluating a large and diverse kidney nuclei dataset, we identified CellViT as having superior performance. However, a performance gap remains between general and kidney-specific nuclei segmentation. Additionally, we utilized all three foundation models to curate consensus “good” and “failure” image patches, summarizing their characteristics to offer insights for improving cell nuclei foundation models in kidney pathology.
6 Acknowledgments

### v2

@@ Abstract
Foundation model, Kidney Pathology, Instance Segmentation, Nuclei, Model Agreement, Ensemble learning
@@ 1 INTRODUCTION
To answer the preceding question, we evaluated the performance of current cell nuclei foundation models in kidney pathology. As shown in Fig. 1, we created a diverse evaluation dataset that includes kidney nuclei data from 2,542 kidney WSIs sourced from humans and rodents across both public and in-house datasets. The dataset includes various tissue types and sizes, such as biopsy, nephrectomy and autopsy samples, and features both formalin-fixed paraffin-embedded (FFPE) and frozen tissues. To our knowledge, the scale of this study’s WSIs surpasses all publicly available labeled nuclei datasets that include the kidney, as illustrated in Fig. 1a. This work conducts a comparative analysis of three widely used SOTA cell nuclei foundation models — Cellpose[8], StarDist [10], and CellViT [9] — on the task of nuclei instance segmentation. For validation, we include a manual rating-based system to curate foundation model predictions and identify failure samples across different models and modalities. To sum up, this study provides a comprehensive assessment of cell nuclei AI foundation models in the field of kidney pathology. The contribution of this paper is three-fold:
$\bullet$ Creation of a Diverse, Large-Scale Kidney Nuclei Dataset: The kidney dataset for evaluation consists of 2,542 kidney WSIs collected from both human and rodent sources, encompassing various tissue types, sizes, and staining methods. To our knowledge, this is the largest-scale evaluation of its kind to date.
$\bullet$ Rating-based segmentation curation with Foundation Models: We utilized three SOTA foundation models (Cellpose[8], StarDist [10], and CellViT [9]) in conjunction with a rating-based system to curate foundation model predictions. Then the evaluation can be shaped as analyzing distribution of the model predictions across the dataset.
$\bullet$ Agreement Analysis between Foundation Models: We performed agreement analysis to identify image patches where all models consistently rated them as “bad.” It allowed us to summarize examples and patterns of these consensus “failure” image patches, offering valuable insights for addressing nuclei segmentation challenges in kidney pathology.
Figure 1: Overall framework. (1) The creation of a diverse, large-scale kidney nuclei dataset.(2) Nuclei instance segmentation with Cellpose [8], StarDist [10], and CellViT [9]. (3) Rating on Segmentation Performance. Meaningful images were then identified through model agreement analysis.
2 METHOD
Fig. 1 illustrates the overall framework of this study: (1) Large-scale diverse dataset creation, (2) Nuclei instance segmentation, (3) Rating on Segmentation Performance. Meaningful images were then identified through model agreement analysis.
@@ 2.2 Nuclei Instance Segmentation
Nuclei instance segmentation was performed on these kidney nuclei image patches using three foundation models. We summarized the three foundation models and provided GitHub links to our customized GPU-optimized inference pipelines in Table. 1:
@@
Year-Month Model Backbone Post-processing Customized GPU Inference
2022 Mar StarDist (Histo.)[10] U-Net Star-convex Polygon Link URL
2022 Nov Cellpose 2.0[8] U-Net GradientFlow Tracking Link URL
2023 Oct CellViT[9] ViT HoVer-Net[15] Link URL
@@ 2.3 Rating on Segmentation Performance
Two pathologist-trained students scored the model predictions as “good,” “medium,” or “bad.” “Good” predictions correctly capture about 90% of nuclei in an image patch, “bad” predictions capture less than 50%, and the rest are categorized as “medium.” We apply this scoring system to quantitatively assess each model’s predictions in our dataset. Finally, we ensemble ratings from three foundation models to curate strongly agreed-upon failure samples. These image patches can be used for training or fine-tuning future foundation models in kidney pathology.
@@ 3 Experiments
3.1 Distribution of Rated Foundation Model Predictions
We analyzed the distribution of predictions for each foundation model across our evaluation dataset. This analysis provides insights into the performance and behavior of each model by examining the frequency of different prediction classes (“good”, “medium”, “bad”) within each foundation model. The prediction class assignments were determined through the manual rating-based curation.
@@ 3.2 Agreement Analysis Between Foundation Models
3.3 Consensus “Bad” Image Patches Mining
@@ 4 RESULTS
4.1 Foundation Model Performance Evaluation
First of all, we rated the segmentation performance of these foundation models. Fig. 2 shows the distribution of rated predictions. CellViT demonstrated superior performance in segmenting nuclei in these kidney images, with 5,609 (63.6%) “good” predictions, which outperforms Cellpose (40.5%) and StarDist (40.2%). This indicates that, despite not being specifically trained on kidney pathology nuclei data, the models retain some transferable knowledge applicable to our task. However, even with the best CellViT model, there is still an existence of approximate 37% of rated predictions from “medium” and “bad,” which underscores the need for a kidney domain-specific foundation model.
@@ 4.1 Foundation Model Performance Evaluation
4.2 Foundation Model Agreement
Fig. 3a illustrates the consistency between model predictions. As shown in the Fig. 3a, no pair exhibits over 90% agreement or complete disagreement. The highest agreement is observed between Cellpose and StarDist (0.76), while the lowest is between Cellpose and CellViT (0.67). These non-extreme values, with a maximum of 0.76 and a minimum of 0.67, indicate that the current SOTA nuclei foundation models in digital pathology do not generalize exceptionally well to diverse large-scale kidney data. Each model has its own strengths, and this inconsistency inspires us to combine multiple SOTA foundation models to enhance performance in downstream kidney nuclei tasks. Both Fig. 3b and Table. 2 show the number of image patches where all three models agree, two models agree, or no models agree for each prediction class (“good”, “medium”, “bad”). It is evident that “bad” cases exhibit the highest inter-model reliability, with over 70% agreement among all three models.
4.3 Ensemble Learning with Foundation Models
Particularly, samples where all three foundation models agree in both the “good” (Consensus “good”) and “bad” (Consensus “failure”) classes are displayed, along with their respective failure categories. As shown, current nuclei foundation models perform well on image patches with darker nuclei and higher contrast between nuclei and non-nuclei regions. Conversely, Consensus “failure” samples exhibit overall lower staining intensity, lighter nuclei, and reduced imaging contrast. Specifically, long and flat nuclei, nuclei with blurred boundaries, and densely distributed nuclei within glomeruli are particularly difficult to segment accurately. Additionally, slides that contain tubules filled with protein casts, or have an excessive number of red blood cells can lead to a higher incidence of false positive predictions.
@@ 4.3 Ensemble Learning with Foundation Models
Table 2: This table shows the percentages of model agreement in each rated prediction class.
@@ 5 Conclusion
This study provides a comprehensive assessment of cell nuclei foundation models (Cellpose, StarDist, CellViT) in kidney pathology, focusing on nuclei instance segmentation. By evaluating a large and diverse kidney nuclei dataset, we identified CellViT as having superior performance. However, a performance gap remains between general and kidney-specific nuclei segmentation. Additionally, we utilized all three foundation models to curate consensus “good” and “failure” image patches, summarizing their characteristics to offer insights for improving cell nuclei foundation models in kidney pathology. An ongoing future direction will further mine these evaluation outcomes to support the training of kidney-domain-specific cell nuclei AI foundation models while minimizing annotation effort.
Acknowledgements.